A man is 24 years older than his son. 12 years ago, he was five times as old as his son.find their present ages

Answer by richwmiller(17219)

m=s+24
m-12=5(s-12)
m=5(s-12)+12
s+24=5(s-12)+12
s+12=5s-60
72=4s
s=18
m=42
check
5*6=30
ok